OpenStudy (anonymous):

use the quadratic formula to find any x-intercepts on the graph of the equation. y=x^2 - 8x -1 x=?

OpenStudy (anonymous):

set y =0, and solve the quadratic equation

OpenStudy (anonymous):

\[x ^{2}-8x-1=0\]

OpenStudy (zehanz):

Set a=1, b=-8 and c=-1 in the QF.

OpenStudy (anonymous):

\[x=4+\sqrt{17} or x=4-\sqrt{17}\]

OpenStudy (anonymous):

yup exactly

OpenStudy (anonymous):

you need to use the quadratic formula

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

so the final answer would be

OpenStudy (anonymous):

\[(4+\sqrt{17}, 0) and (4-\sqrt{17}, 0)\]

OpenStudy (anonymous):

8 square root 68 over 8?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

is it over 4?

OpenStudy (anonymous):

if a =1 yes

OpenStudy (anonymous):

so am i right? it is over 4

OpenStudy (anonymous):

over 2x1